On analysis a substance was found to have the following percentage composition K=28.16,CL=25.63,O=46.21.Determine the molecular formula if it`s molecular mass is 138.5 , write the name of compound

Grade:11

1 Answers

Nishant
6 years ago
            %                Atomicmass         %/mass                  ratio                       simplest form
                                                                                                                                                                 
K –    28.16                39                       28.16/39                0.72                         0.72/0.72    =     1              
Cl-     25.63                35.5                    25.63/35.5            0.72                         0.72/0.72    =      1             
O –    46.21                16                       46.21/16                2.88                         2.88/0.72    =      4
 
                       emperical formula=KClO4
                       emperical  mass=90.5
                         molecular mass= n times emperical  mass
                        138.5=n times 90.5
                             n=138.5/90.5
                             n=1.497
                          molecular formula=(KClo4)3/2